Use formula L=10*log I/Io, where the loudness of a sound in decibels is determined by I, the number of watts per square meter produced by the soundwave, and Io=10^-12 w/m^2. A certain noise produces 6.2 x 10^-4 w/m^2 of power. What is the decibel level of this noise?
Answer by Alan3354(21583)

Use formula L=10*log I/Io, where the loudness of a sound in decibels is determined by I, the number of watts per square meter produced by the soundwave, and Io=10^-12 w/m^2. A certain noise produces 6.2 x 10^-4 w/m^2 of power. What is the decibel level of this noise?
---------------------
L = 10*log(6.2E-4/1E-12)
L = 10*log(6.2E08)
L = 87.9239 dB
L ~ 88 dB